# Detuning-symmetric laser cooling of many mechanical modes with a photothermally modified cavity

### v2

@@ Abstract
Figure 1: Canonical optomechanical system with photothermal effects. a) Schematic. Two mirrors form an optical cavity with total energy decay rate $\kappa=\kappa_{\rm{in}}+\kappa_{\rm{loss}}$. The displacement $x$ of the bottom mirror, a mass-spring harmonic oscillator, modulates the cavity length and detuning $\Delta$ between the cavity resonance (red standing wave) and input laser (top red block) frequencies. Dotted red lines represent heat from absorbed cavity light, which, through thermorefractive and thermoelastic effects, also modulates $\Delta$. b) Feedback loop describing the system, including radiation pressure (black) bolometric forces (blue) and thermal effects in the mirrors (red). For radiation pressure, the nominal mechanical susceptibility ($\chi_{\rm m}$) is modified by a feedback loop comprising optomechanical coupling $G$, the cavity’s susceptibility $C$ to detuning fluctuations $\tilde{\Delta}$ (about mean value $\bar{\Delta}$), and radiation pressure ($-\hbar G$). The red loops describes photothermal mechanisms that modify the bare cavity transfer function, producing the additional stabilization reported here. Here these are dominated by thermoelastic ($\mathcal{T}_{\text{E}}$) and the thermo-refractive ($\mathcal{T}_{\text{R}}$) effects.
Photothermal Model.— As discussed in Ref. [22], photothermal effects offer the possibility of improving mechanical stability via optical spring and optical damping simultaneously. Figure 1(a) shows a schematic of the canonical optomechanical system [1] comprising an optical cavity with one movable mirror acting as a harmonic oscillator of mass $m$, resonant frequency $\Omega_{\text{m}}$, damping rate $\Gamma_{\text{m}}$, and bare mechanical susceptibility
$\chi_{\text{m}}=\frac{1/m}{\Omega_{\text{m}}^{2}-\omega^{2}+i\Gamma_{\text{m}}\omega}.$ (1)
@@
The effects of radiation pressure are well-described by the black feedback loop in Fig. 1(b). To understand how $\chi_{\text{m}}$ changes, we examine how an external oscillatory force $\tilde{F}_{\text{ext}}e^{i\omega t}$ induces a commensurate displacement $\tilde{x}e^{i\omega t}$ (about equilibrium $\bar{x}$). With the loop closed, the total force amplitude $\tilde{F}=\tilde{F}_{\text{ext}}+\tilde{F}_{\text{cav}}$ drives displacement $\tilde{x}=\chi_{\text{m}}\tilde{F}$, which detunes the cavity by $\tilde{\Delta}=G\tilde{x}$ via the optomechanical coupling $G$. This in turn modulates the photon number by $\tilde{n}_{\text{cav}}=\mathcal{C}\tilde{\Delta}$, where
$\mathcal{C}=\left(\frac{\kappa_{\rm in}p_{\text{in}}}{\bar{\Delta}^{2}+\frac{\kappa^{2}}{4}}\right)\left(\frac{1}{\bar{\Delta}+\omega+i\frac{\kappa}{2}}+\frac{1}{\bar{\Delta}-\omega-i\frac{\kappa}{2}}\right),$ (2)
@@
$m\Omega_{G}^{2}$ $\approx\hbar G^{2}\text{Re}\left[\mathcal{C}\right]$ (3)
@@
$\Gamma_{G}$ $\approx-\frac{\hbar G^{2}}{m\Omega_{\text{m}}}\text{Im}\left[\mathcal{C}\right]$ (4)
@@
Importantly, the real (imaginary) part of the cavity’s detuning susceptibility $\mathcal{C}$ determines the optical spring (damping), and both are antisymmetric functions of detuning $\bar{\Delta}$. As such, in the fast-cavity limit ($\kappa>\omega$), stabilizing optical spring (damping) is always accompanied by a destabilizing antidamping (antispring) [1], placing upper bounds on the achievable strength of either.
@@
arising from thermoelastic ($\mathcal{T}_{\text{E}}$) and thermorefractive ($\mathcal{T}_{\text{R}}$) effects, both of which convert cavity photon modulation $\tilde{n}_{\text{cav}}$ directly to detuning modulation $\tilde{\Delta}_{\mathcal{T}}$ by adjusting the effective cavity length. Solving the lower loop equation $\tilde{n}_{\text{cav}}=\tilde{\Delta}\mathcal{C}$ for $n_{\text{cav}}$ alone (now with $\tilde{\Delta}_{\text{m}}$ playing the role of “external drive”) similarly yields an effective cavity detuning susceptibility
$\mathcal{C_{\text{eff}}}=\frac{\mathcal{C}}{1-\mathcal{T}\mathcal{C}}\approx\mathcal{C}+\mathcal{T}\mathcal{C}^{2}$ (6)
when the effect is sufficiently small (e.g., at low power $p_{\text{in}}$ or absorption, when heating is small). The last expression provides quick intuition: In the fast-cavity limit ($\kappa\gg\omega$), $\tilde{n}_{\text{cav}}$ responds nearly adiabatically to $\tilde{\Delta}$, and so $\mathcal{C}$ is mostly real-valued. On the other hand, if thermal time scales are slow compared to mechanical dynamics, the correction term $\mathcal{T}\mathcal{C}^{2}$ can be largely imaginary, providing additional damping or antidamping proportional to $\mathcal{C}^{2}\propto p_{\text{in}}^{2}$ (see Eq. 2), depending on the sign of $\text{Im}[\mathcal{T}]$. Other effects (e.g., dielectric nonlinearities) can also be included in this loop, but these should have a comparatively small effect on detuning, and are anyway too fast to generate a significant imaginary $\mathcal{C}_{\text{eff}}$. In the fast-cavity limit, photothermal effects lead to an additional damping
@@
In practice, thermoelastic effects ($\mathcal{T}_{\text{E}}$) tend to grow mirrors inward and shorten the cavity when power is absorbed, which generates antidamping. However, thermorefractive effects ($\mathcal{T}_{\text{R}}$) tend to lengthen the cavity by increasing the index of refraction in the coatings, thereby generating damping [22].
@@
Fig. 2(a) shows the system used to observe multimode stabilization. The mechanical element is a 3.3 mm $\times$ 3.1 mm, 180 nm thick Si3N4 membrane (patterned into the shown hexagonal lattice) positioned within a 30-$\upmu$m-long fiber optical cavity addressed by two lasers. The first “control” laser (1550 nm) achieves finesse $10^{4}$ (depending on membrane position); the main purpose of this laser is to apply optomechanical forces, but its signal is also used to feedback-stabilize the cavity length. The second “readout” laser (1417 nm) provides a low-finesse ($\lesssim 300$) readout with negligible back-action and sensitivity that is independent of the control laser.
@@
Figure 2: Blue-detuned multimode laser cooling. (a) Schematic. A 1550 nm “control” beam passes through an intensity modulator (IM) and circulator to a fiber cavity near the center of a patterned membrane (inset image). The control beam applies radiation pressure forces via the IM voltage $V_{\text{Drive}}$ and provides the error signal used to stabilize the cavity length via feedback to shear piezos attached to the mirrors. A 1417 nm beam provides essentially backaction-free readout with fixed sensitivity. (b) Displacement spectra recorded by the low-finesse 1417-nm beam. Pink shows the spectrum with the high-finesse beam (1550 nm, 0.5 mW input) blue-detuned to $\bar{\Delta}\sim 3\kappa$, where there is comparatively negligible back-action. Black shows the spectrum for $\bar{\Delta}=\kappa/\sqrt{12}$, where the optical spring is maximal and the membrane’s mechanical noise peaks are suppressed. (c) Pink and black curves show the cumulative variance of the spectra in (b). Blue bars are the ratio of cumulative variances for high and low backaction for smaller ranges of frequencies, while gray bars are simulated (COMSOL) variance ratio expected entirely from changes in mode mass, i.e., in the absence of optical damping. The comparatively low observed values (blue) provide a coarse summary of the optical damping’s frequency dependence.
@@
Figure 1(b) shows two mechanical noise spectra recorded by the readout beam. The black spectrum is taken while the control beam is blue-detuned to $\bar{\Delta}=\kappa/\sqrt{12}$, that is, where the optical spring is strongest. The pink spectrum is recorded at $\bar{\Delta}=3\kappa$ where the optical spring is reduced by a factor of $\sim$70 and the magnitude of the radiation-pressure damping (photothermal servo damping) is reduced by a factor of $\gtrsim$2000 ($\gtrsim$5000). Importantly, in the high-backaction measurement, all prominent peaks – corresponding to the low mass, high quality factor membrane modes – are suppressed: a first hallmark of broadband damping.
@@
Figure 1(c) shows the cumulative variance of the spectra in (b). The low-backaction curve (pink) exhibits pronounced steps at each membrane mode frequency. These are reduced at every step when the backaction is high (black), such that the variance from all modes (rightmost value) is suppressed by a factor of 13 over the full bandwidth. The blue bars show the ratio of integrated variances for the high and low backaction spectra coarsely binned by frequency. In all bins, the remaining variance is more than a factor of 10 smaller. However, some of this suppression is known to arise from the increased stiffness and inertial mass of each mode as it is pinned under the optical spring [25]. To estimate this effect, the gray bins show a COMSOL simulation assuming $\Gamma_{G}=\Gamma_{\mathcal{T}}=0$ to provide an estimate of the suppressed variance expected exclusively from mass and frequency changes. In all cases, the observed variance is reduced by more than this.
@@
Servo modified dynamics— To further validate our interpretation, we directly probe the changes in frequency (Fig. 3(a,b)) and damping (Fig. 3(c,d)) of the fundamental mechanical mode (53.8 kHz) as the high-finesse laser power (at fixed detuning $\bar{\Delta}=\pm\kappa/\sqrt{12}$) and detuning (at fixed incident power 0.5 mW) are varied. The (a) power and (b) detuning dependences of the frequency shift exhibit the linear power dependence and antisymmetric detuning dependence expected for either radiation-pressure dominated or servo-modified backaction, though with a gentle “sag” at higher $\bar{n}_{\text{cav}}$ due to the membrane heating, expanding, and loosening in proportion to intracavity power [25]. However, the damping in (c) exhibits a strong and positive quadratic dependence on power, consistent with photothermal damping (Eq. 7), and qualitatively different from the linear dependence (dashed lines) expected from radiation pressure alone. The detuning dependence in (d) also differs qualitatively from the radiation-pressure expectation, notably exhibiting positive damping for both signs of detuning. The data in Fig. 3 is simultaneously fit (solid lines) to a model incorporating the sum of radiation pressure (Eqs. 3-4), photothermal servo (Eq. 7), and “loosening” due to thermal expansion of the membrane material (modeled simply as a frequency shift $\delta\Omega_{\text{th}}=A/(\bar{\Delta}^{2}/\kappa^{2}+1/4)$ in proportion to cavity power). Dashed lines show the contribution from radiation pressure alone for reference (i.e., $\text{Im}[\mathcal{T}]=0$). This fit reports optomechanical coupling strength $G=1.4$ GHz/nm, thermal loosening coefficient $A=24$ Hz/mW incident, and $\text{Im}[\mathcal{T}]=82$ Hz. The fit indicates optical losses of $(\kappa/2\pi,\kappa_{\rm in}/2\pi)=(3.0,0.2)$ GHz. The agreement of this data with the servo-modified model – along with the striking disagreement with radiation-pressure-force expectations – and the observed simultaneous noise suppression of the first $\gtrsim 100$ modes of a membrane represent the main results of this work.
@@
Figure 3: Power and detuning dependences of dynamical backaction including photothermal effects. (a) Frequency shift and (b) optical damping of the fundamental (53.8 kHz) membrane mode versus incident power. For each power, the laser detuning is set to $\Delta\sim\pm\kappa/\sqrt{12}$, where the optical spring is maximal. Red (blue) dots correspond to laser detunings on the red (blue) side of resonance. Solid curves are simultaneous fits to the entire dataset including thermal loosening and radiation pressure contributions (see main text). (c) Frequency shift and (d) optical damping of the fundamental mode as a function of laser detuning for an incident power $P_{\rm{in}}$=0.5 mW. The dotted lines show the frequency shift and damping expected from radiation pressure alone using parameters extracted from the fit, highlighting the relative contribution of the photothermal servo.
@@
Design considerations.— Whether thermal effects generate positive or negative damping depends on the competition between the thermoelastic expansion (transfer function $\mathcal{T}_{\text{E}}$), which shortens the cavity by growing the mirrors inward, and the thermorefractive index change (transfer function $\mathcal{T}_{\text{R}}$), which effectively lengthens the cavity. Combined with thermal lag at a given frequency $\omega$, the sum of these effects can generate damping only at frequencies $\omega$ where $\mathcal{T}_{\text{R}}$ dominates [26, 20]. As $\omega$ decreases, the thermal penetration depth increases, leading to more material expanding and a larger value of $\mathcal{T}_{\text{E}}$. By contrast, $\mathcal{T}_{\text{R}}$ remains comparatively unchanged, since interference dictates that only the first few layers contain significant electric field. As such, there exists a “crossover” frequency below which thermal expansion dominates, resulting in antidamping. To realize the stability like what is shown here, it is critical to ensure this crossover is below the fundamental frequency of the mechanical system. In previous free-space systems, the crossover frequency was simulated to be $\omega\sim$ 100 kHz for standard mirror coatings, resulting in antidamping of the (lowest-frequency) modes [21]. In our system, the observed damping in Fig. 3 suggests that the crossover frequency is at most the fundamental mode frequency of 53.8 kHz, and $|\mathcal{T}|$ is large enough for the photothermal damping to dominate at modest laser powers. One reason for this is likely our small cavity spot size ($\sim 10~{}\upmu$m diameter), which concentrates the heat deposition to an area $\gtrsim$ 100 times smaller than that of a typical freespace cavity; this leads to faster thermal response, which can enhance the overall magnitude of $|\mathcal{T}|$ for a given frequency (as long as it’s above the inverse thermal time scale [27, 28]).
@@
Figure 4: Enhancement of the thermorefractive effect with modified mirror coatings. (a) Partial removal of the coating’s terminating Ta2O5 layer enhances the electric field in the subsequent SiO2 layer, increasing $\mathcal{T}_{\text{R}}$ by up to a factor of 2.3. Insets show the electric field profile (red) in the cavity (white) and first few layers of a Bragg mirror comprising Ta2O5 (blue) and SiO2 (gray) for 0, 53, 90% removal. (b) A $\lambda/2$ SiO2 spacer embedded in the Bragg stack acts as a low-finesse “mini-cavity”, enhancing the field within it and $\mathcal{T}_{\text{R}}$. Adding more layers after the spacer increases the mini-cavity finesse further boosting $\mathcal{T}_{\text{R}}$, at the expense of added coating thickness. Note for simplicity we assume $\omega$ is low enough that the temperature profile is approximately uniform over the shown layers.
@@
Another promising approach is to design the mirror coating to enhance the field in the SiO2layer(s). Figure 4(b) shows an example of this, where the thickness of one glass layer is doubled, creating a short, low-finesse Fabry-Perot “mini-cavity” within the stack. This concentrates the field in the glass layer, enhancing $\mathcal{T}_{\text{R}}$ as shown in the main plot. Embedding this “mini-cavity” deeper in the stack increases its finesse and enhancement, so long as it is not deeper than thermal penetration depth. On the other hand, if heat is primarily deposited in the mini-cavity itself (where the field is high), then the thermal profile would follow, and very large enhancements are possible. This interplay is the subject of future investigation.

### v3

@@ Abstract
Figure 1: Canonical optomechanical system with photothermal effects. a) Schematic. Two mirrors form an optical cavity with total energy decay rate $\kappa=\kappa_{\rm{in}}+\kappa_{\rm{loss}}$. The displacement $x$ of the bottom mirror, a mass-spring harmonic oscillator, modulates the cavity length and detuning $\Delta$ between the cavity resonance (red standing wave) and input laser (top red block) frequencies. Dotted red lines represent heat from absorbed cavity light, which, through thermorefractive and thermoelastic effects, also modulates $\Delta$. b) Feedback loop describing the system, including radiation pressure (black) bolometric forces (blue) and thermal effects in the mirrors (red). For radiation pressure, the nominal mechanical susceptibility ($\chi_{\rm m}$) is modified by a feedback loop comprising optomechanical coupling $G$, the cavity’s susceptibility $\mathcal{D}$ to detuning fluctuations $\tilde{\Delta}$ (about mean value $\bar{\Delta}$), and radiation pressure ($-\hbar G$). The red loops describes photothermal mechanisms that modify the bare cavity transfer function, producing the additional stabilization reported here. Here these are dominated by thermoelastic ($\mathcal{T}_{\text{E}}$) and the thermo-refractive ($\mathcal{T}_{\text{R}}$) effects.
Photothermal Model.— As discussed in Ref. [22, 24], photothermal effects offer the possibility of improving mechanical stability via optical spring and optical damping simultaneously. Figure 1(a) shows a schematic of the canonical optomechanical system [1] comprising an optical cavity with one movable mirror acting as a harmonic oscillator of mass $m$, resonant frequency $\Omega_{\text{m}}$, damping rate $\Gamma_{\text{m}}$, and bare mechanical susceptibility
$\chi_{\text{m}}=\frac{1/m}{\Omega_{\text{m}}^{2}-\omega^{2}-i\Gamma_{\text{m}}\omega}.$ (1)
@@
The effects of radiation pressure are well-described by the black feedback loop in Fig. 1(b). To understand how $\chi_{\text{m}}$ changes, we examine how an external oscillatory force $\tilde{F}_{\text{ext}}e^{i\omega t}$ induces a commensurate displacement $\tilde{x}e^{i\omega t}$ (about equilibrium $\bar{x}$). With the loop closed, the total force amplitude $\tilde{F}=\tilde{F}_{\text{ext}}+\tilde{F}_{\text{cav}}$ drives displacement $\tilde{x}=\chi_{\text{m}}\tilde{F}$, which detunes the cavity by $\tilde{\Delta}=G\tilde{x}$ via the optomechanical coupling $G$. This in turn modulates the photon number by $\tilde{n}_{\text{cav}}=\mathcal{D}\tilde{\Delta}$, where
$\mathcal{D}=\left(\frac{\kappa_{\rm in}p_{\text{in}}}{\bar{\Delta}^{2}+\frac{\kappa^{2}}{4}}\right)\left(\frac{1}{\bar{\Delta}+\omega+i\frac{\kappa}{2}}+\frac{1}{\bar{\Delta}-\omega-i\frac{\kappa}{2}}\right),$ (2)
@@
$m\Omega_{G}^{2}$ $\approx\hbar G^{2}\text{Re}\left[\mathcal{D}\right]$ (3)
@@
$\Gamma_{G}$ $\approx-\frac{\hbar G^{2}}{m\Omega_{\text{m}}}\text{Im}\left[\mathcal{D}\right]$ (4)
@@
Importantly, the real (imaginary) part of the cavity’s detuning susceptibility $\mathcal{D}$ determines the optical spring (damping), and both are antisymmetric functions of detuning $\bar{\Delta}$. As such, in the fast-cavity limit ($\kappa>\omega$), stabilizing optical spring (damping) is always accompanied by a destabilizing antidamping (antispring) [1], placing upper bounds on the achievable strength of either.
@@
arising from thermoelastic ($\mathcal{T}_{\text{E}}$) and thermorefractive ($\mathcal{T}_{\text{R}}$) effects, both of which convert cavity photon modulation $\tilde{n}_{\text{cav}}$ directly to detuning modulation $\tilde{\Delta}_{\mathcal{T}}$ by adjusting the effective cavity length. Solving the lower loop equation $\tilde{n}_{\text{cav}}=\tilde{\Delta}\mathcal{D}$ for $n_{\text{cav}}$ alone (now with $\tilde{\Delta}_{\text{m}}$ playing the role of “external drive”) similarly yields an effective cavity detuning susceptibility
$\mathcal{D}_{\text{eff}}=\frac{\mathcal{D}}{1-\mathcal{T}\mathcal{D}}\approx\mathcal{D}+\mathcal{T}\mathcal{D}^{2}$ (6)
when the effect is sufficiently small (e.g., at low power $p_{\text{in}}$ or absorption, when heating is small). The last expression provides quick intuition: In the fast-cavity limit ($\kappa\gg\omega$), $\tilde{n}_{\text{cav}}$ responds nearly adiabatically to $\tilde{\Delta}$, and so $\mathcal{D}$ is mostly real-valued. On the other hand, if thermal time scales are slow compared to mechanical dynamics, the correction term $\mathcal{T}\mathcal{D}^{2}$ can be largely imaginary, providing additional damping or antidamping proportional to $\mathcal{D}^{2}\propto p_{\text{in}}^{2}$ (see Eq. 2), depending on the sign of $\text{Im}[\mathcal{T}]$. Other effects (e.g., dielectric nonlinearities) can also be included in this loop, but these should have a comparatively small effect on detuning, and are anyway too fast to generate a significant imaginary $\mathcal{D}_{\text{eff}}$. In the fast-cavity limit, photothermal effects lead to an additional damping
@@
In practice, thermoelastic effects ($\mathcal{T}_{\text{E}}$) tend to grow mirrors inward and shorten the cavity when power is absorbed, which generates antidamping. However, thermorefractive effects ($\mathcal{T}_{\text{R}}$) tend to lengthen the cavity by increasing the index of refraction in the coatings, thereby generating damping [22]. This competition can produce net damping or antidamping depending on the optical, optomechanical, and thermal characteristics of the system.
@@
Fig. 2(a) shows the system used to observe multimode stabilization. The mechanical element is a 3.3 mm $\times$ 3.1 mm, 180 nm thick Si3N4 membrane (patterned into the shown hexagonal lattice) positioned within a 30-$\upmu$m-long fiber optical cavity addressed by two lasers. The first “control” laser (1550 nm) achieves finesse $10^{4}$ (depending on membrane position); the main purpose of this laser is to apply optomechanical forces, but its signal is also used to feedback-stabilize the cavity length via piezoelectric actuators below the fiber mirrors; note the feedback gain is reduced to a value well below the threshold at which it influences the mechanical spectrum (especially the damping parameters). The second “readout” laser (1417 nm) provides a low-finesse ($\lesssim 300$) readout with negligible back-action and sensitivity that is essentially independent of the control laser.
@@
Figure 2: Blue-detuned multimode laser cooling. (a) Schematic. A 1550 nm “control” beam passes through an intensity modulator (IM) and circulator to a fiber cavity near the center of a patterned membrane (inset image). The control beam applies radiation pressure forces via the IM voltage $V_{\text{Drive}}$ and provides the error signal used to stabilize the cavity length via feedback to shear piezos attached to the mirrors. A 1417 nm beam provides a low-finesse, effectively backaction-free readout with fixed sensitivity, calibrated in (b) and (c) against the piezos’ specified response to applied voltage; this provides an approximate reference scale but is not critical to our primary findings, as piezo response uncertainties introduce a common factor to both spectra, preserving their ratio. A wavelength division multiplexer (WDM) separates the two beams. (b) Displacement spectra recorded by the low-finesse 1417-nm beam. Pink shows the spectrum with the high-finesse beam (1550 nm, 0.5 mW input) blue-detuned to $\bar{\Delta}\sim 3\kappa$, where there is comparatively negligible back-action. Black shows the spectrum for $\bar{\Delta}=\kappa/\sqrt{12}$, where the optical spring is maximal and the membrane’s mechanical noise peaks are suppressed. (c) Pink and black curves show the cumulative variance of the spectra in (b). Blue bars are the ratio of cumulative variances for high and low backaction for smaller ranges of frequencies, while gray bars are simulated (COMSOL) variance ratio expected entirely from changes in mode mass, i.e., in the absence of optical damping. The comparatively low observed values (blue) provide a coarse summary of the optical damping’s frequency dependence.
@@
Figure 2(b) shows two mechanical noise spectra recorded by the readout beam. The black spectrum is taken while the control beam is blue-detuned to $\bar{\Delta}=\kappa/\sqrt{12}$, i.e., where the optical spring is strongest and we expect large damping. The pink spectrum is recorded so far detuned ($\bar{\Delta}=3\kappa$) that the optical spring is reduced by a factor of $\sim$70 and the radiation-pressure (photothermal servo) damping by a factor $\gtrsim$2000 ($\gtrsim$5000). This approach allows us to closely approximate “turning off” the optomechanics without sacrificing the robust, consistent lock achieved with the (low-noise) control beam. Importantly, in the high-backaction measurement, all prominent peaks – corresponding to the low mass, high quality factor membrane modes – are suppressed: a first hallmark of broadband damping.
@@
Figure 2(c) shows the cumulative variance of the spectra in (b). The low-backaction curve (pink) exhibits pronounced steps at each membrane mode frequency. These are reduced at every step when the backaction is high (black), such that the variance from all modes (rightmost value) is suppressed by a factor of 13 over the full bandwidth. The blue bars show the ratio of integrated variances of the two spectra (including all modes) coarsely binned by frequency. In all bins, the remaining variance is more than a factor of 10 smaller. However, some of this suppression is known to arise from the increased stiffness and inertial mass of each mode as it is pinned under the optical spring [26]. To estimate this effect, the gray bins show a COMSOL simulation assuming $\Gamma_{G}=\Gamma_{\mathcal{T}}=0$ to provide an estimate of the suppressed variance expected exclusively from mass and frequency changes. In all cases, the observed variance is reduced by more than this, indicating broadband damping.
@@
Servo modified dynamics— To clearly distinguish between ‘force-like’ dynamics [24] (black and blue loops in Fig. 1(b)) and ‘servo-like’ (red loop in Fig. 1(b)), we directly probe the changes in frequency (Fig. 3(a,b)) and damping (Fig. 3(c,d)) of the fundamental mechanical mode (53.8 kHz) as the control beam’s power (at fixed detuning $\bar{\Delta}=\pm\kappa/\sqrt{12}$) and detuning (at fixed incident power 0.5 mW) are varied. The (a) power and (b) detuning dependences of the frequency shift exhibit the linear power dependence and antisymmetric detuning dependence expected for either radiation-pressure dominated or servo-modified backaction, though with a gentle “sag” at higher $\bar{n}_{\text{cav}}$ due to the membrane heating, expanding, and loosening in proportion to intracavity power [26] (other modes follow similar behavior). However, the damping in (c) exhibits a strong and positive quadratic dependence on power, consistent with photothermal damping (Eq. 7), and qualitatively different from the linear dependence (dashed lines) expected from radiation pressure alone. The detuning dependence in (d) also differs qualitatively from the radiation-pressure expectation, notably exhibiting positive damping for both signs of detuning. The data in Fig. 3 is simultaneously fit (solid lines) to a model incorporating the sum of radiation pressure (Eqs. 3-4), photothermal servo (Eq. 7), and loosening due to thermal expansion of the membrane material (modeled simply as a frequency shift $\delta\Omega_{\text{th}}=A/(\bar{\Delta}^{2}/\kappa^{2}+1/4)$ in proportion to cavity power). Dashed lines show the contribution from radiation pressure alone for reference (i.e., $\text{Im}[\mathcal{T}]=0$). This fit reports optomechanical coupling strength $G=1.4$ GHz/nm, thermal loosening coefficient $A=24$ Hz/mW incident, and $\text{Im}[\mathcal{T}]=82$ Hz. The fit indicates optical losses of $(\kappa/2\pi,\kappa_{\rm in}/2\pi)=(3.0,0.2)$ GHz. The agreement of this data with the servo-modified model – along with the striking disagreement with radiation-pressure-force expectations – and the observed simultaneous noise suppression of the first $\gtrsim 100$ modes of a membrane represent the main results of this work.
@@
Figure 3: Power and detuning dependences of dynamical backaction including photothermal effects. (a) Frequency shift and (c) optical damping of the fundamental (53.8 kHz) membrane mode versus incident power. For each power, the laser detuning is set to $\Delta\sim\pm\kappa/\sqrt{12}$, where the optical spring is maximal. Red (blue) dots correspond to laser detunings on the red (blue) side of resonance. Solid curves are simultaneous fits to the entire dataset including thermal loosening and radiation pressure contributions (see main text). (b) Frequency shift and (d) optical damping of the fundamental mode as a function of laser detuning for an incident power $P_{\rm{in}}$=0.5 mW. The dotted lines show the frequency shift and damping expected from radiation pressure alone using parameters extracted from the fit, highlighting the relative contribution of the photothermal servo.
@@
Design considerations.— Whether thermal effects generate positive or negative damping depends on the competition between the thermoelastic expansion (transfer function $\mathcal{T}_{\text{E}}$), which shortens the cavity by growing the mirrors inward, and the thermorefractive index change (transfer function $\mathcal{T}_{\text{R}}$), which effectively lengthens the cavity. Combined with thermal lag at a given frequency $\omega$, the sum of these effects can generate damping only at frequencies $\omega$ where $\mathcal{T}_{\text{R}}$ dominates [27, 20]. As $\omega$ decreases, the thermal penetration depth increases, leading to more material expanding and a larger value of $\mathcal{T}_{\text{E}}$. By contrast, $\mathcal{T}_{\text{R}}$ remains comparatively unchanged, since interference dictates that only the first few layers contain significant electric field. As such, there exists a “crossover” frequency below which thermal expansion dominates, resulting in antidamping. To realize the stability like what is shown here, it is critical to ensure this crossover is below the fundamental frequency of the mechanical system. In previous free-space systems, the crossover frequency was simulated to be $\omega/2\pi\sim$ 16 kHz for standard mirror coatings, resulting in antidamping of the (lowest-frequency) modes [21]. In our system, the observed damping in Fig. 3 suggests that the crossover frequency is at most the fundamental mode frequency of 53.8 kHz, and $|\mathcal{T}|$ is large enough for the photothermal damping to dominate at modest laser powers. One reason for this is likely our small cavity spot size ($\sim 10~{}\upmu$m diameter), which concentrates the heat deposition to an area $\gtrsim$ 100 times smaller than that of a typical freespace cavity; this leads to faster thermal response, which can enhance the overall magnitude of $|\mathcal{T}|$ for a given frequency (as long as it’s above the inverse thermal time scale [28, 29]).
@@
Figure 4: Enhancement of the thermorefractive effect with modified mirror coatings. (a) Partial removal of the coating’s terminating Ta2O5 layer enhances the electric field in the subsequent SiO2 layer, increasing $\mathcal{T}_{\text{R}}$ by up to a factor of 2.3. Insets show the electric field profile (red) in the cavity (white) and first few layers of a Bragg mirror comprising Ta2O5 (blue) and SiO2 (gray) for 0, 53, 90% removal. (b) A half-wavelength SiO2 spacer embedded in the Bragg stack acts as a low-finesse “mini-cavity”, enhancing the field within it and $\mathcal{T}_{\text{R}}$. Adding more layers after the spacer increases the mini-cavity finesse further boosting $\mathcal{T}_{\text{R}}$, at the expense of added coating thickness. Note for simplicity we assume $\omega$ is low enough that the temperature profile is approximately uniform over the shown layers.
@@
Another promising approach is to design the mirror coating to enhance the field in the SiO2layer(s). Figure 4(b) shows an example of this, where the thickness of one glass “spacer” layer is doubled, creating a half-wavelength, low-finesse Fabry-Perot “mini-cavity” within the stack. This concentrates the field in the glass layer, enhancing $\mathcal{T}_{\text{R}}$ as shown in the main plot. Embedding this “mini-cavity” deeper in the stack increases its finesse – and hence the enhancement shown in Fig. 4(b), so long as it is not deeper than thermal penetration depth. On the other hand, if heat is primarily deposited in the mini-cavity itself (where the field is high), then the thermal profile would follow, and very large enhancements are possible. Even in this case, however, the mini-cavity should not be embedded too deeply for two reasons. First, since the mini-cavity is (by design) resonant with the cavity light, it acts to transmit the incident cavity light through approximately twice the depth of the spacer layer (i.e., the mini-cavity comprises a comparable number of layers on both sides of the spacer), thereby requiring additional dielectric layers to maintain the finesse of the main cavity. Second, a mini-cavity of effective length $l$ and finesse $\mathcal{F}$ will localize light outside the main cavity (see inset figures), thereby reducing the optical coupling to the object(s) of interest in the main cavity. Intuitively, whenever circulating photons pass through a mini-cavity, they travel an extra effective path length $\sim l\mathcal{F}$, thereby increasing the overall mode volume. This effect will be small so long as the main cavity length $L\gg l\mathcal{F}$, which should be straightforward for the shallow cavities shown in Fig. 4(b).